OpenStudy (anonymous):

which is the equation of a line that passes through the point (-10 3) and is perpendicular to y=5x-7

OpenStudy (anonymous):

y=5x+53 y=- (1/5)x-7 y=-(1/5)x+1 y=(1/5)x+5

OpenStudy (anonymous):

I think it is b, am I correct?

OpenStudy (anonymous):

nvm then